OpenStudy (devan):

can someone help me find the slope for y= -5/3x + 8/3

OpenStudy (anonymous):

its -5/3. If it is in the form y=ax+b, a is always going to be your slope.

OpenStudy (devan):

so is the answer -5/3 ?

OpenStudy (anonymous):

yes.

OpenStudy (kainui):

At my school we learned y=mx+b which is exactly the same thing. Just remember that m is the slope of your line (rise over run) and b is your y-intercept (where the line goes across y when x=0)

OpenStudy (devan):

so how would i do this equation: y+6=-1/6(x+3) ?

OpenStudy (kainui):

Get it into y=mx+b form. Start out by distributing -1/6 across the (x+3) to get -1/6x - 1/2 and then subtract 6 from both sides to get y=-1/6x-13/2